OpenStudy (anonymous):

Which equation shows a vertical stretch of 2 ,a reflection over the x-axis,then a shift left 3 units and down 4 units.

OpenStudy (anonymous):

A.y=-2/x+4-3 B.y=-2/x+3-4 C.y=-(2/x+4-3) D.y=-(2/x+3-4)